### v1

# Experimental Measurements of the Granular Density of Modes via Impact

Abstract
The jamming transition is an important feature of granular materials, with prior work showing an excess of low frequency modes in the granular analog to the density of states, the granular density of modes. In this work, we present an experimental method for acoustically measuring the granular density of modes using a single impact event to excite vibrational modes in an experimental, three dimensional, granular material. We test three different granular materials, all of which are composed of spherical beads. The first two systems are monodisperse collections of either 6 mm or 8 mm diameter beads. The third system is a bidisperse mixture of the previous two bead sizes. During data collection, the particles are confined to a box; on top of this box, and resting on the granular material, is a light, rigid sheet onto which pressure can be applied to the system. To excite the material, a steel impactor ball is dropped on top of the system. The response of the granular material to the impact pulse is recorded by piezoelectric sensors buried throughout the material, and the density of modes is computed from the spectrum of the velocity autocorrelation of these sensors. Our measurements of the density of modes show more low frequency modes at low pressure, consistent with previous experimental and numerical results, as well as several low frequency peaks in the density of modes that shift with applied pressure. Finally, we also see that the density of modes falls off at a wavelength on the order of twice the particle diameter, which is reminiscent of the Debye frequency.
I Introduction
Ranging from the rings of Saturn to a pile of sand on the beach, granular materials are ubiquitous throughout the Universe, yet many important granular phenomena are not fully understood. One such phenomena is jamming [1], in which a granular material can undergo a transition from solid-like to liquid-like behavior or vice versa. Jamming is an important phenomena for understanding not only the fundamental behavior of granular materials but also many industrial [16] and geophysical [5, 17] processes. For example, how close or far a material is from the jamming transition is studied in geophysics for the purposes of predicting soil stability and events such as landslides. Additionally, jamming also has application for the manufacturing of pharmaceuticals as much of this process involves the manipulation of granular materials [11].
One particularly interesting phenomena that is associated with the jamming transition is that an excess number of vibration modes (soft modes) are seen as the material approaches the jamming transition. As such, the density of states (also known as the density of modes when applied to granular materials) can be used to study the jamming transition, and potentially be used to characterize how far the material is from the jamming transition. Several previous simulations and experiments have done this for granular [8, 14, 18, 13, 4, 2] and colloidal systems [3, 6] using a variety of methods. However, many of the current methods used to measure the granular density of modes are cumbersome and not well suited to use outside of a laboratory setting. This work seeks to develop a simplified measurement process, particularly in regards to phonon excitation, that will aid in more widespread accessibility and deployment of density of modes measurements in non-laboratory, field conditions.
In this paper, we will experimentally measure the density of modes of a real three-dimensional granular material utilizing the spectrum of the velocity autocorrelation function. This is a standard method used in solid state physics [7] and has been experimentally applied to granular materials [13, 4, 2]. A brief summary of the method is as follows: first, the velocity autocorrelation function, $C_{v}(T)$, of a thermal system is given by
$C_{v}(t)=\frac{\sum_{j}\langle v_{j}(\tau+t)\cdot v_{j}(\tau)\rangle_{\tau}}{\sum_{j}\langle v_{j}(\tau)\cdot v_{j}(\tau)\rangle_{\tau}}$ (1)
where $v_{j}(t)$ is the velocity as a function of time of the $j$ particle, and the sum is over all the particles in the system. Then, the density of states, $D(\omega)$, is given by the Fourier transform of the velocity autocorrelation function
$D(\omega)=\int_{0}^{\infty}C_{v}(t)e^{-i\omega t}\mathrm{d}t$ (2)
In particular, when using the normalized autocorrelation function defined in Equation 1, the density of modes is normalized as follows:
$\int_{0}^{\omega_{D}}D(\omega)\mathrm{d}\omega=1$ (3)
where $\omega_{D}$ is the highest frequency mode in the system, and in the Debye model, $\omega_{D}$ has a wavelength that is twice the inter-atomic separation.
This method of finding the density of states assumes that the system is in thermal equilibrium and that the equipartition theorem holds. It also assumes access to the velocities of all the particles. However, even though granular materials are athermal systems, this method was applied to a 2D granular material by exciting the granular material with a white noise acoustic signal to mimic thermal motion [13]. This complex method of exciting the granular material proved successful and agreed well with previous simulations, even though it only utilized a subset of the particles. Building on this method, another experiment [4] used the natural phonon emissions of stick slip events of a granular material under shear to compute the density of modes. A key aspect of utilizing the spectrum of $C_{v}$ to find the granular $D(\omega)$ is that the granular material needs to be perturbed so that velocities can be measured.
In this paper, we present a novel approach to exciting the granular material so that the granular density of modes can be found using the spectrum of the velocity autocorrelation function. Our method excites the vibrational modes using a small impact on top of the granular material. This impact will create a pulse containing a wide spectrum of frequencies. The range of frequencies in a pulse is inversely proportional to the width of the pulse, with an infinitesimal pulse having a flat spectrum. This wide injection of frequencies will allow the granular material to vibrate more strongly at its natural resonances and attenuate signals that are not resonant with the natural modes of the system. This will effectively produce quasi-thermal vibrations among the grains.
This method is functionally very similar to the white noise excitation used in [13]; however, generating an impact is experimentally simple, leading to an improved technique that can be more broadly applied. Our results show that this technique yields a density of states that is consistent with the prior work in that the system has an excess of low frequency modes as the jamming transition is approached. Additionally, we are able to identify several peaks in the density of modes that shift with pressure (distance from the jamming transition).
Figure 1: (a) A schematic of the experimental setup. The granular material, shown as the dotted fill, is confined inside a box, shown as the black line surrounding the granular material. Inside the granular material, labeled with a “P” in the figure are the piezoelectric sensors. On top of the granular material is the “Pressure Plate” assembly. The pressure plate assembly houses the force sensors that measure the pressure applied to the granular material via the “Screw” which is itself confined to the “Frame.” Vibrations are excited in the granular material when the “Ball” is released from the solenoid and dropped through a tube onto the bottom plate of the pressure plate assembly. The bottom plate of the pressure plate assembly is in direct contact with the top of the granular material. (b) A photo showing the ball dropping apparatus with the red wires composing the solenoid and the aluminum tube below the solenoid. To the right, and from the top down, are as follows: the piezoelectric sensor encased in the ABS half spheres, the bare piezoelectric sensor, a 6 mm grain, a 8 mm grain, and the steel impactor ball.
II Experimental Methods
[…]
Figure 2: A typical charge versus time signal from one of the piezoelectric sensors. The multiple pulses are due to the ball bouncing on the pressure plate; the time between the first and second bounce is approximately 0.1 seconds. The inset is a zoomed in view of the first pulse.
II.1 Phonon Excitation
Next, since the granular material is athermal and not in continuous motion, it must be excited in order to vibrate. In order to excite the granular material to vibrate at its natural modes, we drop a 1.05 g, 6.35 mm diameter, steel ball, see Figure 1, on top of the system, thereby injecting an acoustic pulse into the system. As described earlier, a pulse is used due to the fact that it contains a wide range of frequencies, and will produce vibrations in the granular material similar to thermal vibrations.
In order to produce a consistent pulse across our different packings and pressures, the following apparatus was constructed to drop the pulse-producing steel ball on the granular material. The pulse producing apparatus (labeled “ball” in Figure 1) consists of a 5.0 cm tall hollow aluminum tube on top of which is a solenoid. Inside and at the top of the tube, but below the solenoid, the steel ball is placed. The steel ball is held in place by the solenoid’s magnetic field and can be dropped when the solenoid is turned off. This setup allows for a consistent kinetic energy of the ball (0.5 mJ) as it strikes the top plate of the system, as well as provides a means of coordinating the time at which the ball is dropped with the data acquisition of the embedded piezoelectric sensors. Additionally, by striking the rigid plate of the pressure plate apparatus, the energy of the pulse is more evenly distributed across the granular material rather than being concentrated in a particular region of the system.
II.2 Piezoelectric Sensors
We use five piezoelectric sensors embedded in the granular material to measure the acoustic vibrations of the system caused by the steel ball being dropped on the granular material. The particular piezoelectric sensors used are Murata 7BB-12-9 buzzers. While these crystals are marketed as buzzers, they work equally well as detectors due to the symmetric nature of the piezoelectric effect: that is, a strain on a piezoelectric crystal will produce charge separation and, inversely, a charge difference across a piezoelectric crystal will produce a strain. The piezoelectric sensors are 12 mm in total diameter with the crystal being 9 mm in diameter. For each sensor, we glue two ABS plastic half spheres on either side in order to turn the sensor into a 12 mm diameter ABS sphere with a piezoelectric sensor running through its equator, see Figure 1. Encasing the sensors in ABS spheres allows the sensor to better blend in with the rest of the material which is also composed of ABS plastic spheres. This helps to ensure that the sensors are vibrating in as close a manner to the bulk granular material as possible.
[…]
Figure 3: A conceptual schematic of the piezoelectric pre-processing circuit (details are omitted). $A_{1}$ and $A_{2}$ together with $R_{f}$ and $C_{f}$ form the two charge amplifiers and set the low frequency response of the system. The two resistors, $R_{i}$, together with the piezoelectric’s capacitance set the high frequency response. Finally, $A_{3}$ is the differential amplifier that gives the final output voltage proportional to the charge difference on the piezoelectric.
Figure 4: A histogram of the “time of flight” sound speeds for the 6 mm monodisperse packing. Each of the 20 speed measurements was conducted on a unique packing and with no confining pressure. The average sound speed was found to be $193\pm 16$ m/s (standard error).
If these sensors are simply placed into the granular material without any additional circuitry, a voltage could be measured in some proportion to acoustic vibrations. However, this signal would be very weak, possibly not even exceeding the noise threshold. It is therefore necessary to connect the piezoelectric sensors to pre-processing circuitry to amplify the signal and reduce the noise level. The circuit that we developed (shown in Figure 3) to accomplish this task consists of two charge amplifiers [10], one connected to each side of the crystal. These charge amplifiers convert the charge on each side of the crystal into a voltage. These two voltages are then fed through a differential amplifier for both amplification and common mode noise rejection. This circuit yields high sensitivity and low noise, ideal for measuring small vibrations. Additionally, the output voltage of the circuit, $V_{out}$ can be converted into the charge, $q$, generated by the piezoelectric with the following relation $q=\frac{V_{out}}{2GC_{f}}$, where $G$ is the gain of the differential amplifier and $C_{f}$ is the feedback capacitor.
This circuit also introduces low and high frequency cut offs for our measurements. The low frequency cutoff is set by the feedback resistor and capacitor in the charge amplifiers. For our circuit values, this gives a 16Hz, -3dB frequency. A high frequency cut off is intentionally introduced in order to roll-off high frequency noise that is outside the frequency range of interest, and is set by the input resistance $R_{i}$ and the capacitance of the piezoelectric crystal (8 nF). Using our circuit values gives a -3dB frequency of 100kHz; however, in practice, since our sample rate is 50 kHz, we only measure up to 25 kHz due to the Nyquist frequency.
II.3 Data Collection Protocol
The data collection procedure is described as follows: First, all the grains are removed from the box, and mixed in a separate container before being poured back into the experimental apparatus. Removing and replacing the grains ensures that we put the granular material into a new, unique state each time we collect data. Then, once the granular material has been placed back into the experimental apparatus, we place the pressure plate on top of the grains, but do not apply any pressure with the screw. At this point, we drop the steel ball onto the system exciting the vibrational modes. At the same time, the solenoid is de-energized and the steel ball is released, the DAQ card is triggered and begins to record the signal from the piezoelectric sensors for one second. Figure 2 shows a typical output of one of the piezoelectric sensors, with the multiple pulses corresponding to the steel ball bouncing on the pressure plate.
After this, pressure is applied via the screw to the pressure plate, and the ball is dropped again and the response of the piezoelectrics is again recorded. This procedure is repeated for a total of four pressure states (listed above). Once all the pressure states have been cycled through. The grains are removed, mixed, added back into the experimental apparatus, and the whole procedure is repeated. In total, we perform this procedure on five unique packings for each of our three grain types (15 unique packings in total).
Figure 5: The density of modes for (a) the bidisperse system, (b) the 8 mm monodisperse system, and (c) the 6 mm monodisperse system. The different colors correspond to the different pressures that the system is subjected to.
Figure 5: The density of modes for (a) the bidisperse system, (b) the 8 mm monodisperse system, and (c) the 6 mm monodisperse system. The different colors correspond to the different pressures that the system is subjected to.
Figure 6: (a) The DOM for the bidisperse packing zoomed in on the frequencies between 10 Hz and 1 kHz. The location of the primary peaks are labeled with a plus sign and the secondary peak with a star symbol. (b) The frequency of the primary peak with the inset showing the corresponding wavelengths for the 6 mm beads. (c) The frequency of the secondary peak with the inset showing the corresponding wavelengths for the 6 mm beads.
II.4 Sound Speeds
In order to be able to relate specific frequencies to relevant length scales, we need to know the speed of sound in our granular material. This was accomplished for the 6 mm packing by measuring the time difference between the first arrival of sound (time of flight) at two piezoelectric sensors separated from each other by 10 cm. For these sound speed measurements, one piezoelectric sensor is placed near the top of the system and the other is placed 10 cm below this. A pulse of sound is injected at the top of the system and the time at which the signal first arrives at each sensor is found. From this time of flight information, a sound speed for the material is found. The grains are then agitated to put the material in a new state and the sound speed measurement is repeated for a total of 20 unique grain configurations. A histogram of these speeds is shown in Figure 4. We find an average speed of $193\pm 16$ m/s (standard error).
III Results and Analysis
Using the piezoelectric charge as a function of time data (see Figure 2 for an example), we can calculate the granular density of modes via the spectrum of the velocity autocorreclation function. As a first step towards calculating the density of modes, we claim that the the average velocities of the grains around a sensor particle is proportional to the charge on the piezoelectric crystal, that is $q\propto v$ where $q$ is the charge on the crystal and $v$ is the average velocity of the grains around the sensor. Our reasoning for this conclusion is as follows.
First, we recognize that the charge generated on the piezoelectric is proportional to the strain on the crystal, that is $q\propto x$, where $x$ is the change in thickness of the crystal. From this, we can see from both an electrical and mechanical perspective that the energy imparted to the crystal from the granular material is proportional to $q^{2}$. The electrical perspective treats the piezoelectric crystal as a capacitor, and the energy of any capacitor proportional to $q^{2}$. The mechanical perspective treats the crystal as a spring with the energy of a spring proportional to $x^{2}$ which for a piezoelectric is also proportional to $q^{2}$.
Second, since the piezoelectric sensors are only in contact with the granular material, all of the energy imparted to the piezoelectric must come from the kinetic energy of its neighboring grains. Since the kinetic energy of the grains is proportional to $v^{2}$ and the energy of the piezoectric crystal is proportional to $q^{2}$, it follows that $v\propto q$ where the constant of proportionality is some combination of material properties of the grains and the crystal, for example Young’s modulus, geometry, piezoelectric charge coefficient, etc. The fact that $q$ does not have the precise conversion factor to exactly find $v$ in proper units is not necessary for computing the DOM since Equation 1 simply needs to know the relative amplitudes of $v$ at each frequency, and then the normalization of the DOM will happen automatically as a consequence of Equation 1.
Figure 7: The density of modes for the lowest system pressure for all three grain types.
Now that the velocity of the particles around the sensors can be deduced, the rest of the calculation follows straightforwardly. For each particle type, i.e. 6 mm monodisperse, 8 mm monodisperse, or bidisperse, and at each pressure the velocity autocorrelation function, $C_{v}(t)$, is found by finding the average charge autocorrelation function for all the sensors. Then, as described in Equation 2, the FFT of $C_{v}$ gives the density of modes. Now that we have obtained the density of modes, a final smoothing step is taken by applying a median filter with a 15 Hz window to the density of modes. Compared to the 25 kHz data range, a 15 Hz filter window does not substantially change the density of modes other than a slight flattening of some of the peaks, but it does smooth some of the noise fluctuations seen in the data. The results of these calculations are shown Figure 6. These results are reported in the frequency range of 10 Hz to 25 kHz. 10 Hz is the lowest frequency that we report due to the fact that the impactor ball has a bouncing period of approximately 10 Hz and the low frequency 3dB point of our piezoelectric circuitry is 16 Hz. The high frequency cutoff of 25 kHz is set by the Nyquist frequency.
Some features that are immediately apparent from the density of modes in Figure 6 are the excess low frequency modes compared to the Debye scaling as well as a drop off in the number of modes around 10 kHz. We also see several peaks in the approximately 10–1000 Hz range; for a better visual example, Figure 5(a) is a zoomed in graph of the density of modes in this frequency range. Looking at Figure 5(a), there are two peaks of particular interest that shift with system pressure. The first peak is the overall maximum peak in the system, call this the primary peak, the second peak is around 100 Hz, call this the secondary peak, which also shifts with system pressure. The change in the frequency, $\Delta f$, at which each of these peaks occurs relative to the low pressure peak is found and plotted as a function of system pressure in Figures 5(b) and 5(c), and it is seen that the peaks move to higher frequency with increasing system pressure.
In order to explore these two peaks in the density of modes, we calculate the wavelength corresponding to these frequencies using the sound speed and the 6 mm beads. However, as was shown in [12] the sound speed in a granular material is not constant with respect to pressure, and for spheres following the Hertzian contact law [9], the sound speed should increase with pressure as $P^{1/6}$ [12]. Applying the $P^{1/6}$ correction to the sound speed, we calculate the wavelengths corresponding to the primary and secondary peaks for the 6 mm beads and show the results in the inset in Figures 5(b), and 5(c).
Finally, we also plot the density of modes at the lowest pressure for all three grain types on the same plot in Figure 7 to compare the effects of grain size and compare monodisperse to bidisperse packings.
IV Discussion
One of the first things seen in Figure 6 is that around 10 kHz the density of modes falls off and then levels off to a minimum value. In the Debye model of the density of states, the highest phonon mode has a wavelength of twice the inter-atomic spacing. If we attempt to translate the concept of inter-atomic spacing to a granular system, we might expect the shortest wavelength to be twice the grain diameter. Using the sound speed for our 6 mm monodisperse system leads to a frequency of $16\pm 1.3$ kHz which is within a factor of two from the experimental value. We also see that at the lowest pressure, the density of modes reaches its high frequency cutoff at a lower frequency than when the system is under higher pressure.
The next observation is that the primary peak, shown in Figures 5(a) and 5(b), increases with increasing pressure, and the corresponding wavelength for the 6 mm beads ranges from approximately 50 cm to 60 cm (see inset in Figure 5(b)). These wavelengths are approximately twice the system width and may correspond to a standing wave of the system size, which would explain its dominance in the density of modes. An interesting method to validate this hypothesis would be to create a system that had different dimensions for the length, width, and height of the box and see if multiple peaks in the density of modes emerge.
The secondary peak in Figure 5(c) is also seen to increase with system pressure, and the corresponding wavelengths of the 6 mm beads are shown in the inset in Figure 5(c). The wavelengths of these peaks corresponds to approximately 2 m which is much larger than the system size. The wavelengths also seem to increase with pressure in contrast to the behavior of the primary peak. The fact that these peaks do change with system pressure gives us confidence that they are real features of the granular material and not spurious electrical noise or artifacts. Additionally, there are several other peaks that also appear to scale with system pressure around the 200 Hz mark. These peaks are also soft modes of the system and deserve future investigation.
Another peak of note in the density of modes is the peak seen at 60 Hz. This peak is slightly flattened due to the median filter and the logarithmic scaling of the figure but is well centered at 60 Hz regardless of system pressure. This peak is almost certainly an electrical artifact of the 60 Hz mains in the building’s electrical wiring.
An additional point of investigation in this study is the effect of different grain size as well as monodisperse packings versus bidisperse packings. Looking at Figure 7, we see that while there are some differences in the density of modes between the three systems, it is not a dramatic difference, with general agreement across the main features of the density of modes. This is not particularly surprising given that the grains are all made of the same material and of the same shape (spherical). Additionally, while [13] did see a difference between monodisperse packings and bidisperse packing in 2D due to the crystallization of the 2D packing, no such regular crystallization occurs in 3D.
While there are broad similarities in the density of modes across the different packings, they are not identical. This can be seen in Figures 5(b) and 5(c). In particular, in both plots there is a difference between the 6 mm and 8 mm packings. Curiously, in Figure 5(b) the bidisperse packing closely resembles the 8 mm packing, and then in Figure 5(c) the bidisperse packing switches and closely resembles the 6 mm packing.
Finally, one of the strengths of this work is the novel approach to exciting the vibrational modes with a simple impact event. In addition to being experimentally simple, an impact has a particular intrinsic advantage over other excitation methods. Namely, while the impactor is striking the granular material, the system can be thought of as a damped, driven oscillator, and since the impact event is very short, transients will be important to the response. For damped, driven oscillators, the transient terms oscillate at the natural frequency of the system and not the driving frequency, which is exactly what is needed for our experiments. Then, after the impact event, the driving force is no longer present and the system behaves more like a damped oscillator, essentially ringing at the natural frequencies of the material.
To fully take advantage of this technique it is helpful to make the impact event as short as possible. If the impactor colliding with the impactor plate is thought of as an object colliding with a spring, then the period of that spring’s oscillation would correspond to the duration of the collision. Using this model for the impact time, a short impact can be accomplished with a light, but hard and stiff impactor such as the small, hardened steal ball bearing used in this study. If higher energy is needed from the impactor, it is best accomplished by increasing the speed of the impactor rather than its mass so that the impact time remains small. Finally, so long as the energy of the impactor does not cause the grains to permanently change their position, the energy of the impactor should not influence the $D(\omega)$ of an individual packing, provided there is enough energy to excite the same number of modes with different energies. However, if data from multiple packing are averaged together, as was done in this study, the energy of the impactor should be kept very consistent as packings with higher energy impacts will be weighted more heavily in Equation 1.
V Conclusion
In this work, we have successfully measured the granular density of modes in three dimensions for a packing of spherical grains. Our measurements of the density of modes displays an excess of low frequency modes as the jamming transition is approached. We also identify several interesting peaks that shift with pressure and have length scales at, or longer than, the system size. Future work might vary the dimensions of the granular system in order to see if these peaks shift or split into multiple peaks as the symmetry of the system size is broken. Additionally, we see the density of modes fall off at high frequency in a manner reminiscent of the Debye frequency; that is, the density of states falls off at a wavelength close to twice the particle size. All of these results continue to demonstrate the utility in borrowing solid state physics techniques and applying them to granular systems. Finally, our novel approach of exciting the grains with a simple impact, increases experimental accessibility and will allow for measurements of the granular density of modes to become more commonplace.
Acknowledgements.
We would like to thank the NASA South Carolina Space Grant Consortium and the South Carolina Independent Colleges and Universities for financial support of this work. We would also like to thank Karen Daniels for helpful comments and discussion on this work as well as Thomas Owens for helpful comments and discussions on the experimental apparatus.